OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 20 = x2 − x? (−5, 0) and (4, 0) (5, 0) and (−4, 0) (−5, 0) and (−4, 0) (5, 0) and (4, 0)

OpenStudy (anonymous):

Always for finding x intercept, put y = 0.. \[0 + 20 = x^2 - x\] Subtract 20 both the sides: \[x^2 - x - 20 = 0\] Now factorize it to find x..

OpenStudy (anonymous):

is it the third option

OpenStudy (anonymous):

Show me that what you got after factorization??

OpenStudy (anonymous):

ok give me a sec i sorta did this in a rush

OpenStudy (anonymous):

let me try again see if i get it right this time

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

i got 1/2?

OpenStudy (anonymous):

How ??

OpenStudy (anonymous):

I said to make factors..

OpenStudy (anonymous):

x is aprox 4.944 and x is aprox -3.944

OpenStudy (anonymous):

so then second option?

OpenStudy (anonymous):

Yes second is good but your method is not right..

OpenStudy (anonymous):

oh can you explain

OpenStudy (anonymous):

\[x^2 - x - 20 = 0\] \[x^2 - 5x + 4x - 20 = 0\] \[(x-5)(x+4) = 0\] \[x - 5 = 0 \implies x = 5\] \[x + 4 = 0 \implies x = -4\]

OpenStudy (anonymous):

\[(5,0) \; \; and \; \; (-4, 0)\]

OpenStudy (anonymous):

oh i see what i did wrong i was following the quadratic formula instead of factoring srry